OpenStudy (littlenugget):

Solve the equation graphically by converting it to an equivalent equation with 0 on the right-hand side and then finding the x-intercepts

OpenStudy (littlenugget):

\[2x-5=\sqrt{x+4}\]

OpenStudy (anonymous):

Square both sides.

OpenStudy (littlenugget):

(2x-5)(2x-5)=x+4 4x^2-20x+25=x+4

OpenStudy (anonymous):

subtract x+4 from both sides.

OpenStudy (littlenugget):

4x^2-21x+21

OpenStudy (anonymous):

Now find the roots.

OpenStudy (littlenugget):

so would i factor?

OpenStudy (anonymous):

Yeah, do it anyway you wish.

OpenStudy (littlenugget):